Find the derivative of f(x) = 3x4 - 8x2+2x–7.
Answer #1


Given F (= 3x4–8x*+2X-7 f@)= a C 3x18x*+2x-7) a lode (GM) =nan -(3.41 238.2x+2015-0 plac)= 1243 f (x) 16x+2 =1223 23 168tQ